# TerraCodec: Compressing Optical Earth Observation Data

### v2

@@ Appendix 0.F Qualitative examples
This section complements the main paper with additional visual examples. We first compare reconstructions across all codecs on representative SSL4EO-S12 samples (Sec. 0.F.1). We then examine model beliefs and forecasting behavior (Sec. 0.F.2) for TEC-TT and FlexTEC. We assesss how dropping tokens impacts TEC-TT and the importance of token masking for Latent Repacking. We also show how TEC–TT forecasts the next frame from past context, contrasting mean predictions with stochastic samples.
@@ 0.F.1 General reconstructions
We compare TEC-TT, TEC-FP, and classical codecs (JPEG2000, WebP) on SSL4EO-S12 v1.1 validation samples at matched rate $\approx 0.20$ bppbf. Each row in Fig. 0.F.1 shows the original and reconstructions from each codec, annotated with the average bppbf and PSNR-per-band (65k clipped) across the sequence.
@@ 0.F.2 Model beliefs and forecasting
We discuss the effect of token budget on different TT model versions and show the TEC-TT forecasts.
@@ 0.F.2.1 Token budget comparison
Figure 16: Effect of filling dropped tokens (K=5K{=}5) with model prior predictions at inference for different TEC-$T$ variants. Vanilla TEC-TT exhibits spatial holes and banding when tokens are removed. Adding Latent Repacking without masking improves quality but leaves uneven detail, while FlexTEC preserves global layout and reduces artifacts.
We visualize token budget effects in Fig. 16 using two example sequences under an aggressive token limit, illustrating how models behave when later tokens are dropped. We compare TEC-TT, TEC-TT with Latent Repacking but no masking, and FlexTEC (with both). FlexTEC degrades smoothly and preserves scene-wide structure, whereas TEC-TT exhibits patch erasure and banding; the Latent Repacking w/o masking variant lies in between, confirming that masking with dynamic rate scaling is essential for stable variable-rate performance.
@@ 0.F.2.2 Forecasts from past context
We probe TEC-TT’s model beliefs by predicting the current frame from past context only, using either the predicted mean or samples from the distribution (Fig. 17). The $\mu$-forecast reliably captures large-scale structure, while sampling from the full prior (mean and variance) expresses context-aware uncertainty, primarily reflecting cloud variability. This illustrates the learned distribution rather than a single point estimate. The conservative single-point forecast ($\mu$-forecast) produces a clear-sky prediction for the next frame, enabling TEC-TT to perform cloud removal as evaluated in AllClear.
@@ Appendix 0.G Additional details on cloud inpainting
We provide extended results for the AllClear cloud inpainting benchmark [66], complementing Sec. 5.3. TEC-TT is applied without task-specific fine-tuning, leveraging only its latent temporal predictions.
@@
We use TEC-TT’s prior mean prediction ($\mu$) as a clear-sky estimate of the next frame, capturing large-scale structure while down-weighting transient noise such as clouds (Sec. 0.F.2.2). Building on this, we adapt TEC-TT for cloud removal by predicting the third input frame from the two previous ones and applying cloud-aware decoding: clear regions retain their original tokens, while cloudy regions are replaced with the transformer’s predicted tokens. Cloud and cloud-shadow masks are available and are downscaled to latent resolution using average pooling.
To provide the clean temporal context, we apply a context reordering heuristic that duplicates the least-cloudy inputs when few clean frames are available. We evaluate 16 variants spanning mask type, mask threshold (0.0 vs. 0.5), context reordering (on vs. off), and decoding policy: Interleave, which predicts only cloudy tokens, and Propagate, which predicts from the first cloudy token onward, plus a pure forecasting baseline that replaces all tokens with predictions. The best setting uses cloud+shadow masks at threshold 0.0, average-pool downscaling, context reordering, and interleave decoding; all cloud-aware variants outperform pure forecasting, which tends to follow seasonal drift rather than reconstruct the target. Intuitively, conditioning only cloudy regions on predictions while preserving clear tokens reduces seasonal bias and allows TEC-TT to focus its temporal prior on reconstructing the true clear-sky target. Throughout the paper (Sec. 5.3 , Sec. 0.G.1), we report results using this best configuration.
Although our introduced codecs are trained on L2A reflectance data, the AllClear benchmark is defined on L1C inputs. To ensure a fair zero-shot evaluation, we therefore pretrained a TEC-TT model on the SSL4EO-S12 v1.1 L1C data modality using a high $\lambda=700$ to focus on reconstruction quality. This model uses the same architecture, hyperparameters, and training procedure as the corresponding L2A variant, and is applied without any task-specific fine-tuning on AllClear.
@@ 0.G.1 Quantitative results
Table 7 reports RMSE, and MAE results for all methods, complementing the PSNR and SSIM results in the main paper zero-shot TEC-TT clearly outperform the heuristics: TEC-TT reaches PSNR $\approx 32.9$ dB and SSIM $\approx 0.917$, compared to LeastCloudy ($30.61$ dB / $0.863$) and Mosaicing ($29.82$ dB / $0.755$). Relative to the strongest prior zero-shot neural method on AllClear (UnCRtainTS, $29.01$ dB / $0.898$ / MAE =0.039=0.039 / RMSE =0.050=0.050), TEC-TT is substantially stronger ($32.86$ dB / $0.917$ / MAE =0.025=0.025 / RMSE =0.034=0.034).
@@ 0.G.2 Qualitative results
Figure 19 presents additional TEC-TT inpainting results on the AllClear test set across varying degrees of cloudiness in the three past input frames. While three context images are shown, TEC-TT uses only the two least cloudy as input. Notably, even under heavily clouded conditions, TEC-TT leverages past context to produce relatively accurate predictions of the target frame.

### v3

@@ Appendix 0.F Qualitative examples
This section complements the main paper with additional visual examples. We first compare reconstructions across all codecs on representative SSL4EO-S12 samples (Sec. 0.F.1). We then examine model beliefs and forecasting behavior (Sec. 0.F.2) for TEC-$T$ and FlexTEC. We assesss how dropping tokens impacts TEC-$T$ and the importance of token masking for Latent Repacking. We also show how TEC–$T$ forecasts the next frame from past context, contrasting mean predictions with stochastic samples.
@@ 0.F.1 General reconstructions
We compare TEC-$T$, TEC-FP, and classical codecs (JPEG2000, WebP) on SSL4EO-S12 v1.1 validation samples at matched rate $\approx 0.20$ bppbf. Each row in Fig. 0.F.1 shows the original and reconstructions from each codec, annotated with the average bppbf and PSNR-per-band (65k clipped) across the sequence.
@@ 0.F.2 Model beliefs and forecasting
We discuss the effect of token budget on different $T$ model versions and show the TEC-$T$ forecasts.
@@ 0.F.2.1 Token budget comparison
Figure 16: Effect of filling dropped tokens (K=5K{=}5) with model prior predictions at inference for different TEC-$T$ variants. Vanilla TEC-$T$ exhibits spatial holes and banding when tokens are removed. Adding Latent Repacking without masking improves quality but leaves uneven detail, while FlexTEC preserves global layout and reduces artifacts.
We visualize token budget effects in Fig. 16 using two example sequences under an aggressive token limit, illustrating how models behave when later tokens are dropped. We compare TEC-$T$, TEC-$T$ with Latent Repacking but no masking, and FlexTEC (with both). FlexTEC degrades smoothly and preserves scene-wide structure, whereas TEC-$T$ exhibits patch erasure and banding; the Latent Repacking w/o masking variant lies in between, confirming that masking with dynamic rate scaling is essential for stable variable-rate performance.
@@ 0.F.2.2 Forecasts from past context
We probe TEC-$T$’s model beliefs by predicting the current frame from past context only, using either the predicted mean or samples from the distribution (Fig. 17). The $\mu$-forecast reliably captures large-scale structure, while sampling from the full prior (mean and variance) expresses context-aware uncertainty, primarily reflecting cloud variability. This illustrates the learned distribution rather than a single point estimate. The conservative single-point forecast ($\mu$-forecast) produces a clear-sky prediction for the next frame, enabling TEC-$T$ to perform cloud removal as evaluated in AllClear.
@@ Appendix 0.G Additional details on cloud inpainting
We provide extended results for the AllClear cloud inpainting benchmark [66], complementing Sec. 5.3. TEC-$T$ is applied without task-specific fine-tuning, leveraging only its latent temporal predictions.
@@
We use TEC-$T$’s prior mean prediction ($\mu$) as a clear-sky estimate of the next frame, capturing large-scale structure while down-weighting transient noise such as clouds (Sec. 0.F.2.2). Building on this, we adapt TEC-$T$ for cloud removal by predicting the third input frame from the two previous ones and applying cloud-aware decoding: clear regions retain their original tokens, while cloudy regions are replaced with the transformer’s predicted tokens. Cloud and cloud-shadow masks are available and are downscaled to latent resolution using average pooling.
To provide the clean temporal context, we apply a context reordering heuristic that duplicates the least-cloudy inputs when few clean frames are available. We evaluate 16 variants spanning mask type, mask threshold (0.0 vs. 0.5), context reordering (on vs. off), and decoding policy: Interleave, which predicts only cloudy tokens, and Propagate, which predicts from the first cloudy token onward, plus a pure forecasting baseline that replaces all tokens with predictions. The best setting uses cloud+shadow masks at threshold 0.0, average-pool downscaling, context reordering, and interleave decoding; all cloud-aware variants outperform pure forecasting, which tends to follow seasonal drift rather than reconstruct the target. Intuitively, conditioning only cloudy regions on predictions while preserving clear tokens reduces seasonal bias and allows TEC-$T$ to focus its temporal prior on reconstructing the true clear-sky target. Throughout the paper (Sec. 5.3 , Sec. 0.G.1), we report results using this best configuration.
Although our introduced codecs are trained on L2A reflectance data, the AllClear benchmark is defined on L1C inputs. To ensure a fair zero-shot evaluation, we therefore pretrained a TEC-$T$ model on the SSL4EO-S12 v1.1 L1C data modality using a high $\lambda=700$ to focus on reconstruction quality. This model uses the same architecture, hyperparameters, and training procedure as the corresponding L2A variant, and is applied without any task-specific fine-tuning on AllClear.
@@ 0.G.1 Quantitative results
Table 7 reports RMSE, and MAE results for all methods, complementing the PSNR and SSIM results in the main paper zero-shot TEC-$T$ clearly outperform the heuristics: TEC-$T$ reaches PSNR $\approx 32.9$ dB and SSIM $\approx 0.917$, compared to LeastCloudy ($30.61$ dB / $0.863$) and Mosaicing ($29.82$ dB / $0.755$). Relative to the strongest prior zero-shot neural method on AllClear (UnCRtainTS, $29.01$ dB / $0.898$ / MAE =0.039=0.039 / RMSE =0.050=0.050), TEC-$T$ is substantially stronger ($32.86$ dB / $0.917$ / MAE =0.025=0.025 / RMSE =0.034=0.034).
@@ 0.G.2 Qualitative results
Figure 19 presents additional TEC-$T$ inpainting results on the AllClear test set across varying degrees of cloudiness in the three past input frames. While three context images are shown, TEC-$T$ uses only the two least cloudy as input. Notably, even under heavily clouded conditions, TEC-$T$ leverages past context to produce relatively accurate predictions of the target frame.